OpenStudy (anonymous):

[5.01] Part 1: Solve the inequality 12 + 4x 5x - 2 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (ajprincess):

\(12 + 4x \le 5x - 2\) u hav to isolate x if u r to solve the inequality. To do that follow the steps 1.Subtract 4x both sides. 2.Add 2 both sides. Can u do it @callme911?

OpenStudy (anonymous):

...

OpenStudy (ajprincess):

sorry

OpenStudy (ajprincess):

u dnt get me?

OpenStudy (anonymous):

nop

OpenStudy (ajprincess):

Can u solve this equation?. x+3=5

OpenStudy (anonymous):

x=2

OpenStudy (ajprincess):

yup right:) similarly can u solve 12+4x=5x-2?

OpenStudy (ajprincess):

I am really sorrry. Hav to go nw:(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@callme911 An inequality is practically like an equation. So let's solve it in the same way... 12 + 4x <= 5x - 2 So just like normal equation algebra, you need to get all the x terms on one side and all the number terms on the other. How about subtracting 4x from both sides... 12 + 4x <= 5x - 2 12 + 4x - 4x <= 5x - 4x - 2 12 <= x - 2 Then get all the numbers on the left by adding 2 to both sides... 12 + 2 <= x - 2 + 2 14 <= x That's all there is... it's just like solving an equation. This is what @ajprincess was saying earlier. There is one small difference... doesn't matter here, but if you had an example like -x <= -10 where your x term is negative, you can multiply through by -1 to make it all positive, but when you do that, you have to switch the direction of the inequality: -x <= -10 (-1)(-x) >= (-1)(-10) <<<---- see? "<=" switched to ">=" So.... x >= 10

OpenStudy (anonymous):

what about # 2